# Weighted Fisher divergence for high-dimensional Gaussian variational inference

### v1

@@ 4 Sparse Gaussian variational approximations
Next, we consider Gaussian VI for hierarchical models and compare performances of the Fisher, score-based and KL divergences. Given observed data $y=(y_{1},\dots,y_{n})^{\top}$, the variable $\theta=(\theta_{L}^{\top},\theta_{G}^{\top})^{\top}\in\mathbb{R}^{d}$ of a two-tier hierarchical model can be partitioned into a global variable $\theta_{G}$ that is shared among all observations and local variables $\theta_{L}=(b_{1}^{\top},\dots,b_{n}^{\top})^{\top}$, where $b_{i}$ is specific to the observation $y_{i}$ for $i=1,\dots,n$. Let the joint likelihood of the model be
@@
where $\{y_{i}\}$ are conditionally independent given $\theta$, and the $\{b_{i}\}$ follow an $\ell$th order Markov model given $\theta_{G}$. Thus $\{b_{i}\}$ are conditionally independent of each other a posteriori given the $\ell$ neighboring values and $\theta_{G}$. In a random effects model, $\{b_{i}\}$ are the random effects with $\ell=0$, while for a state space model, $\{b_{i}\}$ are the latent states with $\ell=1$.
Let $q_{\lambda}(\theta)$ be ${\text{N}}(\mu,\Sigma)$, a Gaussian variational approximation of the posterior density with mean $\mu\in\mathbb{R}^{d}$ and covariance matrix $\Sigma\in\mathbb{R}^{d\times d}$. Consider a Cholesky decomposition of the precision matrix $\Omega=\Sigma^{-1}=TT^{\top}$ where $T$ is a lower triangular matrix, and denote the variational parameters as $\lambda=(\mu^{\top},\text{vech}(T)^{\top})^{\top}$, where $\text{vech}(\cdot)$ is an operator that stacks lower triangular elements of a matrix columnwise from left to right into a vector.
In a multivariate Gaussian, conditional independence implies sparse structure in the precision matrix, with $\Omega_{ij}=0$ if $\theta_{i}$ and $\theta_{j}$ are conditionally independent given the remaining variables. By Proposition 1 of Rothman et al. (2010), Cholesky factor $T$ has the same row-banded structure as $\Omega$. Suppose $T$ is block partitioned according to $(b_{1}^{\top},\dots,b_{n}^{\top},\theta_{G}^{\top})^{\top}$, with corresponding blocks $T_{ij}$ for $i,j=1,\dots,n+1$. If $T_{ij}=0$ for $1\leq j\leq i-l$ and $j>i$ (as $T$ is lower-triangular), then $\Omega$ reflects the conditional independence structure of the true posterior. When $\theta$ is high-dimensional, exploiting the conditional independence structure in the model is essential to making Gaussian VI feasible, as the number of parameters to be optimized in $T$ grows quadratically with $n$. However, after imposing sparsity on $T$, the number of parameters only grows linearly with $n$.
We propose two SGD approaches for minimizing the Fisher and score-based divergences, as SGD allows updating of only the elements in $T$ that are not constrained to zero by sparsity. Algorithms that minimize the FD and SD based on unbiased gradient estimates derived using the reparametrization trick are named FDr and SDr (“r” for reparametrization trick, see Table 4), while algorithms that rely on a batch approximation of the FD and SD are named FDb and SDb (“b” for batch approximation, see Table 5) respectively.
@@ 5 SGD based on reparameterization trick
The Fisher and score-based divergences between $q_{\lambda}(\theta)$ and $p(\theta|y)$ can be written as
@@
respectively, where $g(\lambda,\theta)=\nabla_{\theta}\log h(\theta)-\nabla_{\theta}\log q_{\lambda}(\theta)$, $f(\lambda,\theta)=T^{-1}g(\lambda,\theta)$ and $h(\theta)=p(\theta)p(y|\theta)$ as defined previously. We derive gradients of the Fisher and score-based divergences by applying the reparametrization trick of Kingma and Welling (2014). Instead of simulating $\theta$ directly from $q_{\lambda}(\theta)$, we generate $z\sim{\text{N}}(0,I_{d})$ and compute $\theta=\mu+T^{-\top}z$. Thus
@@
where ${\text{E}}_{\phi}(\cdot)$ denotes expectation with respect to $\phi(z)$, the pdf of ${\text{N}}(0,I_{d})$. Note that
@@
Detailed derivations are given in the supplement.
Unbiased estimates of the gradients can be obtained by sampling from $\phi(z)$. All gradient computations can be done efficiently even in high-dimensions, as they only involve sparse matrix multiplications and solutions of sparse triangular linear systems. The Hessian $\nabla_{\theta}^{2}\log h(\theta)$ has the same block sparse structure as $\Omega$, as $b_{i}$ and $b_{j}$ only occur in the same factor of (7) if $b_{j}$ is one of the $\ell$ neighboring values of $b_{i}$. For $\nabla_{\text{\rm vech}(T)}F(\lambda)$ and $\nabla_{\text{\rm vech}(T)}S(\lambda)$, we only need to compute elements corresponding to those in $\text{\rm vech}(T)$ that are not fixed by sparsity. For instance, to compute the second term in $\nabla_{\text{\rm vech}(T)}F(\lambda)$, we just find $u=T^{-\top}z$ and $v=T^{-1}\nabla_{\theta}^{2}h(\theta)g(\lambda,\theta)$, and then form $u_{i}v_{j}$ for nonzero elements $(i,j)$ of $T$.
The update for $T$ in SGD does not ensure positiveness of the diagonal entries. Hence, we introduce $T^{*}$ such that $T^{*}_{ii}=\log(T_{ii})$ for $i=1,\dots,n,$ and $T^{*}_{ij}=T_{ij}\;for\;i\neq j$. Let $J$ be a $d\times d$ matrix with diagonal equal to $\text{\rm diag}(T)$ and all off-diagonal entries being 1, and $D$ be a diagonal matrix with the diagonal given by $\text{\rm vech}(J)$. Then $\nabla_{\text{\rm vech}(T^{*})}F(\lambda)=D\nabla_{\text{\rm vech}(T)}F(\lambda)$ and updates for $T^{*}$ are unconstrained.
KLD FDr and SDr Initialize $\mu$ and $T$. For $t=1,2,\dots N$, 1. Generate $z\sim{\text{N}}(0,I_{d})$. 2. Compute $u=T^{-\top}z$, $\theta=\mu+u$ and $g=\nabla_{\theta}\log h(\theta)+Tz$. 3. $\mu\leftarrow\mu+\rho_{t}g$. 4. $v=T^{-1}g$. 5. $g_{T}=-uv^{\top}$. 6. $T^{*}\leftarrow T^{*}+\rho_{t}\;Dg_{T}$. 7. Obtain $T$ from $T^{*}$. 3. If SD, $g\leftarrow T^{-1}g$. $z\leftarrow z-g$ $g\leftarrow T^{-\top}g$. 4. $w=\nabla_{\theta}^{2}\log h(\theta)g$,
$v=T^{-1}w$.
5. $\mu\leftarrow\mu-2\rho_{t}w$. 6. $g_{T}=gz^{\top}-uv^{\top}$. 7. $T^{*}\leftarrow T^{*}-2\rho_{t}Dg_{T}$. 8. Obtain $T$ from $T^{*}$.
Table 4: SGD algorithms based on reparametrization trick.
Table 4 outlines the SGD algorithms for updating $(\mu,T)$ by minimizing the Fisher, score-based or KL (derived in Tan and Nott, 2018) divergences. The stepsize $\rho_{t}$ is computed elementwise adaptively using Adadelta (Zeiler, 2012). All three algorithms compute $g(\lambda,\theta)$, but the KLD based algorithm uses $g(\lambda,\theta)$ directly to update $\mu$ and $T$, while the FD and SD based algorithms use the Hessian $\nabla_{\theta}^{2}\log h(\theta)$ to condition $g(\lambda,\theta)$ and are hence more computationally intensive.
@@ 5.1 Analysis of variance of gradient estimates
We study variance of the unbiased gradient estimates derived by applying the reparametrization trick on the KL, Fisher and score-based divergences, to examine effects of the conditioning matrix, $\nabla_{\theta}^{2}\log h(\theta)$. The variance of these gradients plays a crucial role in the stability of the optimization, as large variance can cause a zigzag phenomenon, making convergence difficult. For a closed form analysis, we assume the target $p(\theta|y)$ is ${\text{N}}(\nu,\Lambda^{-1})$. Then $\nabla_{\theta}\log h(\theta)=-\Lambda(\theta-\nu)$ and $\nabla_{\theta}^{2}\log h(\theta)=-\Lambda$.
@@
From Table 4, gradient estimates with respect to $\mu$ for the KL, Fisher and score-based divergences based on a single sample are
@@
Similarly, from Table 4, the gradient estimates with respect to $T$ are
@@
The variance of these gradient estimates depends on the mean and precision of the true target, which is fixed, and that of the variational approximation, which keeps changing during SGD. Suppose $\Lambda$ and $T$ are both diagonal matrices, then
@@
It can be verified that these variances are zero at full convergence, when $\mu_{i}=\nu_{i}$ and $T_{ii}^{2}=\Lambda_{ii}$ $\forall$ $i$. The variance of gradients with respect to $\mu$ of FD and SD are larger than that of KLD if $\Lambda_{ii}>0.5$ and $\Lambda_{ii}/T_{ii}^{2}>0.5$ respectively. Assuming $\mu_{i}=\nu_{i}$ for the SD, the variance of gradients with respect to $T$ of FD and SD are larger than that of KLD if $T_{ii}^{2}+\Lambda_{ii}>0.5$ and $\Lambda_{ii}/T_{ii}^{2}>0.25$ respectively.
In summary, the variance of gradient estimates based on FD is larger than that of KL once $\Lambda_{ii}>0.5$ regardless of the values of the variational parameters, and variance inflation is larger for $T$ than $\mu$. For SD, the inflation factor involves the ratio $\Lambda_{ii}/T_{ii}^{2}$, so variance inflation relative to KLD can be reduced if $T_{ii}^{2}>\Lambda_{ii}$.
@@ 6 SGD based on batch approximation
FDb SDb Initialize $\mu$ and $T$. For $t=1,\dots N$, 1. Generate $z_{i}\sim{\text{N}}(0,I_{d})$ and compute $\theta_{i}=\mu+T^{-\top}z_{i}$ for $i=1,\dots B$. 2. Compute $g_{h}(\theta_{i})$ for $i=1,\dots B$ and the summary statistics in (8). 3. $U=C_{\theta}+(\mu-\overline{\theta})(\mu-\overline{\theta})^{\top}$. 4. $g_{\mu}=2TT^{\top}(\mu-\overline{\theta})-2\overline{g}_{h}$. 5. $\mu\leftarrow\mu-\rho_{t}TT^{\top}g_{\mu}$ 6. $W=C_{\theta g}-(\mu-\overline{\theta})\overline{g}_{h}^{\top}$. 7. $g_{T}=2(W+W^{\top}+TT^{\top}U+UTT^{\top})T$. 5. $\mu\leftarrow\mu-\rho_{t}g_{\mu}$. 6. $V=C_{g}+\overline{g}_{h}\overline{g}_{h}^{\top}$. 7. $g_{T}=2(UT-T^{-\top}T^{-1}VT^{-\top})$. 8. $T^{*}\leftarrow T^{*}-\rho_{t}Dg_{T}$. 9. Obtain $T$ from $T^{*}$.
Table 5: SGD algorithms based on batch approximated FD and SD.
Unbiased gradient estimates based on the reparametrization trick involve the Hessian $\nabla_{\theta}^{2}\log h(\theta)$, which is computationally expensive, requires high storage and increases the variability in gradient estimates. This can lead to instability in minimizing the FD and SD. To address these challenges, we propose an alternative approach that eliminates dependence on the Hessian, by minimizing biased estimates of the FD and SD computed using a batch of samples randomly selected from the current variational approximation at each iteration. First, the SD and FD can be written respectively as
@@
where $g_{h}(\theta)=\nabla_{\theta}\log h(\theta)$ and the subscript $q_{\lambda}$ emphasizes that expectation is with respect to $q_{\lambda}(\theta)$. To estimate SD and FD at the $t$-iteration, we generate $B$ samples $\{\theta_{1},\dots,\theta_{B}\}$ from the current estimate of the variational density $q_{t}(\theta)={\text{N}}(\theta|\mu^{(t)},\Sigma^{(t)})$. This can be done by generating $z_{i}\sim{\text{N}}(0,I_{d})$ and computing $\theta_{i}=\mu^{(t)}+T^{(t)-\top}z_{i}$ for $i=1,\dots,B$, where $\Sigma^{(t)}=T^{(t)-\top}T^{(t)-1}$. Let
$\begin{gathered}\overline{\theta}=\frac{1}{B}\sum_{i=1}^{B}\theta_{i},\quad\overline{g}_{h}=\frac{1}{B}\sum_{i=1}^{B}g_{h}(\theta_{i}),\quad\\ C_{\theta g}=\frac{1}{B}\sum_{i=1}^{B}(\theta_{i}-\overline{\theta})(g_{h}(\theta_{i})-\overline{g}_{h})^{\top},\\ C_{\theta}=\frac{1}{B}\sum_{i=1}^{B}(\theta_{i}-\overline{\theta})(\theta_{i}-\overline{\theta})^{\top},\quad\\ C_{g}=\frac{1}{B}\sum_{i=1}^{B}(g_{h}(\theta_{i})-\overline{g}_{h})(g_{h}(\theta_{i})-\overline{g}_{h})^{\top},\end{gathered}$ (8)
be summary statistics based on the batch of samples. Estimates of SD and FD are
$\hat{S}_{q_{t}}(\lambda)$ $=\frac{1}{B}\sum_{i=1}^{B}\left\{g_{h}(\theta_{i})^{\top}\Sigma g_{h}(\theta_{i})+2g_{h}(\theta_{i})^{\top}(\theta_{i}-\mu)\right.$
@@
where $U=C_{\theta}+(\mu-\overline{\theta})(\mu-\overline{\theta})^{\top}$, $V=C_{g}+\overline{g}_{h}\overline{g}_{h}^{\top}$ and $W=C_{\theta g}-(\mu-\overline{\theta})\overline{g}_{h}^{\top}$. Differentiating with respect to $\mu$ and $T$,
@@
These gradient estimates of SD and FD are biased because the $\theta$’s are replaced by samples $\{\theta_{1},\dots,\theta_{B}\}$ generated from $q_{t}(\theta)={\text{N}}(\theta|\mu^{(t)},\Sigma^{(t)})$, and are no longer functions of $(\mu,\Sigma)$ when we derive the gradients. The dependence on $\lambda$ through $\theta$ has been cut. On the other hand, the reparametrization trick in Section 5 produces unbiased estimates because the $\theta$’s are regarded as samples from $q(\theta)={\text{N}}(\theta|\mu,\Sigma)$, and remain functions of $(\mu,\Sigma)$ when the chain rule is applied to find the gradients.
With the batch approximation, all gradients are independent of the Hessian, which reduces computation costs significantly and enhances stability during optimization. As before, we only update elements of $\text{\rm vech}(T)$ not fixed by sparsity, and ensure positivity of diagonal entries in $T$ by applying a transformation. SGD algorithms for updating $(\mu,T)$ based on minimizing the batch approximated FD and SD are outlined in Table 5.

### v2

@@ 4 Sparse Gaussian variational approximations
Next, we consider Gaussian VI for hierarchical models and compare performances of the FD, SD and KLD. Given observed data $y=(y_{1},\dots,y_{n})^{\top}$, the variable $\theta=(\theta_{L}^{\top},\theta_{G}^{\top})^{\top}\in\mathbb{R}^{d}$ of a two-tier hierarchical model can be partitioned into a global variable $\theta_{G}$ that is shared among all observations and local variables $\theta_{L}=(b_{1}^{\top},\dots,b_{n}^{\top})^{\top}$, where $b_{i}$ is specific to the observation $y_{i}$ for $i=1,\dots,n$. Let the joint likelihood of the model be
@@
[…]
Let $q_{\lambda}(\theta)$ be ${\text{N}}(\mu,\Sigma)$, a Gaussian variational approximation of the posterior with mean $\mu\in\mathbb{R}^{d}$ and covariance matrix $\Sigma\in\mathbb{R}^{d\times d}$. Consider a Cholesky decomposition of the precision matrix $\Omega=\Sigma^{-1}=TT^{\top}$ where $T$ is a lower triangular matrix, and denote the variational parameters as $\lambda=(\mu^{\top},\text{vech}(T)^{\top})^{\top}$, where $\text{vech}(\cdot)$ is an operator that stacks lower triangular elements of a matrix columnwise from left to right into a vector.
In a multivariate Gaussian, conditional independence implies sparse structure in the precision matrix, with $\Omega_{ij}=0$ if $\theta_{i}$ and $\theta_{j}$ are conditionally independent given the remaining variables. By Proposition 1 of Rothman et al. (2010), the Cholesky factor $T$ has the same row-banded structure as $\Omega$. Suppose $T$ is block partitioned according to $(b_{1}^{\top},\dots,b_{n}^{\top},\theta_{G}^{\top})^{\top}$, with corresponding blocks $T_{ij}$ for $i,j=1,\dots,n+1$. First, $T_{ij}=0$ if $j>i$ as $T$ is lower-triangular. If we further constrain $T_{ij}=0$ for $1\leq j\leq i-l$, then $\Omega$ reflects the conditional independence structure of the joint likelihood in (7). For instance, for GLMMs with $\ell=0$, $T$ has the sparse block structure,
$T=\begin{bmatrix}T_{11}&0&\ldots&\ldots&0\\ 0&T_{22}&\ldots&\ldots&0\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ 0&0&\dots&T_{nn}&0\\ T_{G1}&T_{G2}&\ldots&T_{Gn}&T_{GG}\\ \end{bmatrix}$.
When $\theta$ is high-dimensional, exploiting the conditional independence structure in the model is essential to making Gaussian VI feasible, as the number of parameters to be optimized in $T$ grows quadratically with $n$. However, after imposing sparsity on $T$, the number of parameters only grows linearly with $n$. Predetermined sparsity in $T$ can be enforced in SGD by updating only the elements in $T$ that are not constrained to zero.
@@ 5 SGD based on reparameterization trick
In this section, we develop SGD algorithms to minimize the FD and SD based on unbiased gradient estimates derived using the reparameterization trick (Kingma and Welling, 2014), named FDr and SDr respectively. In this approach, the gradients involve Hessians of the log joint density, which are sparse matrices that can be computed efficiently. However, we demonstrate later in Section 5.1 that these gradients have much higher variance than corresponding algorithms based on KLD, resulting in slow convergence and suboptimal variational approximations. An alternative approach is thus proposed in Section 6.
Let $g(\lambda,\theta)=\nabla_{\theta}\log h(\theta)-\nabla_{\theta}\log q_{\lambda}(\theta)$ where $h(\theta)=p(\theta)p(y|\theta)$ is as defined previously. The FD and SD between $q_{\lambda}(\theta)$ and $p(\theta|y)$ can be written as
@@
respectively, where $f(\lambda,\theta)=T^{-1}g(\lambda,\theta)$. The gradients for minimizing the FD and SD via SGD can be derived by applying the reparametrization trick. Instead of simulating $\theta$ directly from $q_{\lambda}(\theta)$, we generate $z\sim{\text{N}}(0,I_{d})$ and compute $\theta=\mu+T^{-\top}z$. Thus
@@
where ${\text{E}}_{\phi}(\cdot)$ denotes expectation with respect to $\phi(z)$, the density function of ${\text{N}}(0,I_{d})$. Note that
@@
Unbiased gradient estimates can be obtained by sampling from $\phi(z)$. All gradient computations can be done efficiently even in high-dimensions, as they only involve sparse matrix multiplications and solutions of sparse triangular linear systems. The Hessian $\nabla_{\theta}^{2}\log h(\theta)$ has the same block sparse structure as $\Omega$, as $b_{i}$ and $b_{j}$ only occur in the same factor of (7) if $b_{j}$ is one of the $\ell$ neighboring values of $b_{i}$. For $\nabla_{\text{\rm vech}(T)}F(\lambda)$ and $\nabla_{\text{\rm vech}(T)}S(\lambda)$, we only need to compute elements corresponding to those in $\text{\rm vech}(T)$ that are not fixed by sparsity. For instance, to compute the second term in $\nabla_{\text{\rm vech}(T)}F(\lambda)$, we just find $u=T^{-\top}z$ and $v=T^{-1}\nabla_{\theta}^{2}h(\theta)g(\lambda,\theta)$, and then form $u_{i}v_{j}$ for nonzero elements $(i,j)$ of $T$.
The update for $T$ in SGD does not ensure that its diagonal entries remain positive. Hence, we introduce $T^{*}$ such that $T^{*}_{ii}=\log(T_{ii})$ for $i=1,\dots,n,$ and $T^{*}_{ij}=T_{ij}\;for\;i\neq j$. Let $J$ be a $d\times d$ matrix with diagonal equal to $\text{\rm diag}(T)$ and all off-diagonal entries being 1, and $D$ be a diagonal matrix with the diagonal given by $\text{\rm vech}(J)$. Then $\nabla_{\text{\rm vech}(T^{*})}F(\lambda)=D\nabla_{\text{\rm vech}(T)}F(\lambda)$ and updates for $T^{*}$ are unconstrained.
Algorithm 1 SGD based on reparametrization trick
1:Initial $\mu\in\mathbb{R}^{d}$, initial $T^{*}\in\mathbb{R}^{d\times d}$, stepsize schedule $\{\rho_{t}\}$
2:function Map($T^{*}$)
3:  Construct $T$: $T_{ii}\leftarrow\exp(T^{*}_{ii})$, $T_{ij}\leftarrow T^{*}_{ij}$ for $i\neq j$
4:  return $T$
5:end function
6:function BuildD($T$)
7:  $J\leftarrow\mathbf{1}\mathbf{1}^{\top}$, set $\mathrm{diag}(J)\leftarrow\mathrm{diag}(T)$
8:  $D\leftarrow\mathrm{diag}(\text{\rm vech}(J))$
9:  return $D$
10:end function
$1t\leftarrow 1$
12:while not converged do
13:  $T\leftarrow\textsc{Map}(T^{*})$, $D\leftarrow\textsc{BuildD}(T)$
14:  Sample $z\sim\mathcal{N}(0,I_{d})$, $u\leftarrow T^{-\top}z$, $\theta\leftarrow\mu+u$
15:  $g\leftarrow\nabla_{\theta}\log h(\theta)+Tz$
16:  if KLD then
17:   $\mu\leftarrow\mu+\rho_{t}\,g$, $v\leftarrow T^{-1}g$, $g_{T}\leftarrow-\,u\,v^{\top}$
18:   $T^{*}\leftarrow T^{*}+\rho_{t}\,D\,g_{T}$
19:  else if FDr or SDr then
20:   if SDr then
21:     $g\leftarrow T^{-1}g$, $z\leftarrow z-g$, $g\leftarrow T^{-\top}g$
22:   end if
23:   $w\leftarrow\nabla_{\theta}^{2}\log h(\theta)\,g$, $v\leftarrow T^{-1}w$, $\mu\leftarrow\mu-2\,\rho_{t}\,w$
24:   $g_{T}\leftarrow g\,z^{\top}-u\,v^{\top}$, $T^{*}\leftarrow T^{*}-2\,\rho_{t}\,D\,g_{T}$
25:  end if
26:  $t\leftarrow t+1$
27:end while
Algorithm 1 outlines the SGD algorithms for updating $(\mu,T)$ by minimizing the FD, SD or KLD (derived in Tan and Nott, 2018). The stepsize $\rho_{t}$ is computed elementwise adaptively using Adadelta (Zeiler, 2012). All three algorithms compute $g(\lambda,\theta)$, but the KLD based algorithm uses $g(\lambda,\theta)$ to update $\mu$ and $T$ directly, while FDr and SDr premultiply $g(\lambda,\theta)$ by the Hessian $\nabla_{\theta}^{2}\log h(\theta)$ and are hence more computationally intensive.
@@ 5.1 Analysis of variance of gradient estimates
Here, we study the variance of unbiased gradient estimates derived by applying the reparametrization trick on the KLD, FD and SD. The variance of these gradients plays a crucial role in stability of the optimization, as large variance can cause a zigzag phenomenon, making convergence difficult. For a closed form analysis, we assume the target $p(\theta|y)$ is ${\text{N}}(\nu,\Lambda^{-1})$. Then $\nabla_{\theta}\log h(\theta)=-\Lambda(\theta-\nu)$ and $\nabla_{\theta}^{2}\log h(\theta)=-\Lambda$.
@@
From Algorithm 1, gradient estimates with respect to $\mu$ for the KLD, FD and SD based on a single sample are
@@
Similarly, from Algorithm 1, the gradient estimates with respect to $T$ are
@@
The variance of these estimates depends on the mean and precision of the true target, which is fixed, and that of the variational approximation, which changes during SGD. Suppose $\Lambda$ and $T$ are both diagonal matrices, then
@@
It can be verified that these variances are zero at convergence, when $\mu_{i}=\nu_{i}$ and $T_{ii}^{2}=\Lambda_{ii}$ $\forall$ ii. The variance of gradients with respect to $\mu$ of FD and SD are larger than that of KLD if $\Lambda_{ii}>0.5$ and $\Lambda_{ii}/T_{ii}^{2}>0.5$ respectively. Assuming $\mu_{i}=\nu_{i}$ for the SD, the variance of gradients with respect to $T$ of FD and SD are larger than that of KLD if $T_{ii}^{2}+\Lambda_{ii}>0.5$ and $\Lambda_{ii}/T_{ii}^{2}>0.25$ respectively.
In summary, the variance of gradient estimates based on FD is larger than that of KLD once $\Lambda_{ii}>0.5$, regardless of the values of the variational parameters, and variance inflation is larger for $T$ than $\mu$. For SD, the inflation factor involves the ratio $\Lambda_{ii}/T_{ii}^{2}$, so variance inflation relative to KLD can be reduced if $T_{ii}^{2}>\Lambda_{ii}$.
@@ 6 SGD based on batch approximation
SGD based on the reparametrization trick faces multiple issues such as increased computational and storage costs due to the Hessian, and high variance in gradient estimates. To address these challenges, we propose alternative algorithms (named FDb and SDb) in this section, which minimize estimates of the FD and SD computed using a batch of samples randomly simulated from the current variational approximation at each iteration. In this approach, the gradients are biased but they no longer depend on the Hessian, leading to reduced computation costs and improved convergence. In Section 6.1, we discuss how this approach iteratively refines the variational approximation by emulating gradients of the target posterior evaluated on each batch of samples. This approach is also more scalable and stable compared to BaM (Cai et al., 2024) for high-dimensional models with conditional independence structure that can be exploited. Section 6.2 analyzes the behavior of FDb and SDb under the mean-field assumption in the limit of an infinite batchsize.
Algorithm 2 SGD based on batch approximation
1:Initial $\mu\in\mathbb{R}^{d}$, initial $T^{*}$, batchsize $B$, stepsize schedule $\{\rho_{t}\}$
2:while not converged do
3:  $T\leftarrow\textsc{Map}(T^{*})$, $D\leftarrow\textsc{BuildD}(T)$
4:  Sample $z_{i}\sim\mathcal{N}(0,I_{d})$
5:  $\theta_{i}\leftarrow\mu+T^{-\top}z_{i}$ for $i=1,\dots,B$
6:  Compute $g_{h}(\theta_{i})$ for $i=1,\dots,B$
7:  Compute summary statistics:
$\overline{\theta}$ $\leftarrow\tfrac{1}{B}\sum_{i=1}^{B}\theta_{i},\quad C_{\theta g}\leftarrow\tfrac{1}{B}\sum_{i=1}^{B}(\theta_{i}-\overline{\theta})(g_{h}(\theta_{i})-\overline{g}_{h})^{\top},$
$\overline{g}_{h}$ $\leftarrow\tfrac{1}{B}\sum_{i=1}^{B}g_{h}(\theta_{i}),\quad C_{\theta}\leftarrow\tfrac{1}{B}\sum_{i=1}^{B}(\theta_{i}-\overline{\theta})(\theta_{i}-\overline{\theta})^{\top},$
$C_{g}$ $\leftarrow\tfrac{1}{B}\sum_{i=1}^{B}(g_{h}(\theta_{i})-\overline{g}_{h})(g_{h}(\theta_{i})-\overline{g}_{h})^{\top}$
8:  $U\leftarrow C_{\theta}+(\mu-\overline{\theta})(\mu-\overline{\theta})^{\top}$, $g_{\mu}\leftarrow 2TT^{\top}(\mu-\overline{\theta})-2\overline{g}_{h}$
9:  if FDb then
10:   $\mu\leftarrow\mu-\rho_{t}TT^{\top}g_{\mu}$, $W\leftarrow C_{\theta g}-(\mu-\overline{\theta})\overline{g}_{h}^{\top}$
11:   $g_{T}\leftarrow 2\big(W+W^{\top}+TT^{\top}U+UTT^{\top}\big)T$
12:  else if SDb then
13:   $\mu\leftarrow\mu-\rho_{t}g_{\mu}$, $V\leftarrow C_{g}+\overline{g}_{h}\overline{g}_{h}^{\top}$
14:   $g_{T}\leftarrow 2\big(UT-T^{-\top}T^{-1}VT^{-\top}\big)$
15:  end if
16:  $T^{*}\leftarrow T^{*}-\rho_{t}Dg_{T}$
17:  $t\leftarrow t+1$
18:end while
The SD and FD can be written respectively as
@@
where $g_{h}(\theta)=\nabla_{\theta}\log h(\theta)$ and the subscript $q_{\lambda}$ emphasizes that expectation is with respect to $q_{\lambda}(\theta)$. To estimate SD and FD at the $t$-iteration, we generate $B$ samples $\{\theta_{1},\dots,\theta_{B}\}$ from the current estimate of the variational density $q_{t}(\theta)={\text{N}}(\theta|\mu^{(t)},\Sigma^{(t)})$. This can be done by generating $z_{i}\sim{\text{N}}(0,I_{d})$ and computing $\theta_{i}=\mu^{(t)}+T^{(t)-\top}z_{i}$ for $i=1,\dots,B$, where $\Sigma^{(t)}=T^{(t)-\top}T^{(t)-1}$. By using the summary statistics computed in step 6 of Algorithm 2, estimates of SD and FD at iteration $t$ are
$\hat{S}_{q_{t}}(\lambda)$ $=\frac{1}{B}\sum_{i=1}^{B}\left\{g_{h}(\theta_{i})^{\top}\Sigma g_{h}(\theta_{i})+2g_{h}(\theta_{i})^{\top}(\theta_{i}-\mu)\right$. (8)
@@
where $U=C_{\theta}+(\mu-\overline{\theta})(\mu-\overline{\theta})^{\top}$, $V=C_{g}+\overline{g}_{h}\overline{g}_{h}^{\top}$, $W=C_{\theta g}-(\mu-\overline{\theta})\overline{g}_{h}^{\top}$ and the subscript $q_{t}$ indicates that samples are drawn from $q_{t}$. Differentiating with respect to $\mu$ and $T$,
@@
These gradient estimates of SD and FD are biased because the $\theta$’s are replaced by samples $\{\theta_{1},\dots,\theta_{B}\}$ generated from $q_{t}(\theta)={\text{N}}(\theta|\mu^{(t)},\Sigma^{(t)})$, and are no longer functions of $(\mu,\Sigma)$ when we derive the gradients. On the other hand, the reparametrization trick in Section 5 produces unbiased estimates because the $\theta$’s are regarded as samples from $q(\theta)={\text{N}}(\theta|\mu,\Sigma)$, and remain functions of $(\mu,\Sigma)$ when the chain rule is applied to find the gradients.
With the batch approximation, all gradients are independent of the Hessian, which reduces computation costs significantly and enhances stability during optimization. As before, we only update elements of $\text{\rm vech}(T)$ not fixed by sparsity, and ensure positivity of diagonal entries in $T$ by applying a transformation. SGD algorithms for updating $(\mu,T)$ based on minimizing the batch approximated FD and SD are outlined in Algorithm 2.